OpenStudy (anonymous):

Find the area of the region between y=x^(1/2) and y=x^(1/3) for 0≤x≤1

OpenStudy (loser66):

can you draw them out?

OpenStudy (anonymous):

go to here and type in the functions https://www.desmos.com/calculator

OpenStudy (anonymous):

Im interested in understanding the concepts behind the problem- I have the answer (.0833) This is in calculus, using definite integrals.

OpenStudy (anonymous):

No graph is included in the question.

OpenStudy (anonymous):

So the formula is \[\int\limits_{a}^{b} TOP function - BOTTOM function\]

OpenStudy (loser66):

@VostheBoss321 but we have to know which one is upper function in interval.

OpenStudy (anonymous):

f(b)-f(a)?

OpenStudy (anonymous):

You can graph it yourself and you probably know what x^1/2 looks like

OpenStudy (loser66):

I drew it, take a look, and follow ConDawg https://www.desmos.com/calculator

OpenStudy (anonymous):

I know what both graphs look like, but is the only way to sole this problem graphically? Or is there an analytical method?

OpenStudy (anonymous):

You need to graph it first it help to understand things a lot better and it doesnt take long.

OpenStudy (loser66):

you see,in your interval, 0,1, upper is x^1/3 , out of the interval, upper is x^1/2. No way to solve without graph

OpenStudy (anonymous):

go to that website and type in y=x^1/2 y=x^1/3 it takes two seconds

OpenStudy (anonymous):

Alright, I'm looking at the graph from site Condawg provided, and I see the space in question

OpenStudy (anonymous):

How do I compute the area of the region though?

OpenStudy (anonymous):

\[\int\limits_{0}^{1}x ^{1/3}-x ^{1/2}\]

OpenStudy (anonymous):

That simple? Just integral f(b)-f(a)?

OpenStudy (anonymous):

integral of the top function (x^1/3) minus the integral of the bottom function (x^1/2)

OpenStudy (anonymous):

Piece of cake, eh?

OpenStudy (anonymous):

Yup! Thank you both very much!

OpenStudy (anonymous):

tell me what you get for an answer though!

OpenStudy (anonymous):

What value should I be using for x in this case? Or do I apply power rule? (1 and 0 raised to any power end up canceling each other out in above expression)

OpenStudy (anonymous):

Do you know how to take the integral of x^1/2 and x^1/3?

OpenStudy (anonymous):

Frankly, ive been at this a few hours now and I feel theyve slipped my mind

OpenStudy (anonymous):

I do understand how to take integrals though

OpenStudy (anonymous):

Don't worry I'll show you!

OpenStudy (anonymous):

in this case, is it 1/1.5x^1.5 and 1/1.3x^1.3?

OpenStudy (anonymous):

+c of course

OpenStudy (anonymous):

This is the first step is to write it like this! Write this down on your paper! \[\int\limits_{0}^{1}x ^{1/3}-\int\limits_{0}^{1}x ^{1/2}\]

OpenStudy (anonymous):

P.S. use don't use "c" for definite integrals, but that's irrelevant.

OpenStudy (anonymous):

I understand, thats implicit in the whole definite integral concept, correct?

OpenStudy (anonymous):

Yeah, but I'm going to go through this step by step, and you write each step down. So write step one down.

OpenStudy (anonymous):

Done

OpenStudy (anonymous):

Step two: evaluate the integrals To evaluate these integrals we are going to use a method I call the anti derivative method. This method is add 1 to the exponent and divide the whole thing by the new exponent. So: \[\frac{ 3 }{ 4 }*x ^{\frac{ 4 }{ 3 }}0 \to 1\] and \[\frac{ 2 }{ 3 }x ^{\frac{ 3 }{ 2 }} 0 \to 1\]

OpenStudy (anonymous):

Questions?

OpenStudy (anonymous):

I need to look at the anti-derivative method again, we have learned it but in a later chapter.

OpenStudy (anonymous):

But I understand your logic herein completely.

OpenStudy (anonymous):

Step 3: evaluate both the integrals from 0 to 1 then subtract the top function by the bottom function to evaluate the 3/4*x^4/3 we plug in 1 then evaluate then subtract that by what we get by plugging 0 and evaluating. So: \[([\frac{ 3 }{ 4 }(1)^{\frac{ 4 }{ 3 }}) - (\frac{ 3 }{ 4 }(0)^{\frac{ 4 }{ 3 }})]\] that's for the top function the bottom function: \[\frac{ 2 }{ 3 }(1)^{\frac{ 3 }{ 2 }} - \frac{ 2 }{ 3 }(0)^{\frac{ 3 }{ 2 }}\] the top function simplifies to \[\frac{ 3 }{ 4 } - 0\] The bottom function simplifies to \[\frac{ 2 }{ 3 } - 0 = \frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

Alright, thats clear

OpenStudy (anonymous):

last step!

OpenStudy (anonymous):

I got the answer

OpenStudy (anonymous):

step 4: subtract what we evaluated for the top function from the bottom function. So: top= 3/4 bottom= 2/3 \[\frac{ 3 }{ 4 } - \frac{ 2 }{ 3 } = \frac{ 9 }{ 12 } - \frac{ 8 }{ 12 } = \frac{ 1 }{ 12 }\] 1/12

OpenStudy (anonymous):

(3/4)-(2/3)

OpenStudy (anonymous):

right

OpenStudy (anonymous):

or .0833....

OpenStudy (anonymous):

use fraction for integration, it's easier to work with, and in university you can't use calculators for calculus

OpenStudy (anonymous):

So, its essentially just taking the antiderivative, then inputing top and bottom values, simplifying said values, and subtracting top value from bottom

OpenStudy (anonymous):

You got her right there!

OpenStudy (anonymous):

Thank you kindly! As an aside, my professor (I am in college, taking calculus 1 and 2 over the summer to avoid scheduling conflicts) is allowing calculators. Is there another method involving them?

OpenStudy (anonymous):

Nope. You just gotta follow those steps. I'm on here almost every day so feel free to ask me for help.

OpenStudy (anonymous):

remember to graph.

OpenStudy (anonymous):

Thank you very much, i'll be sure to take you up on the offer. And absolutely- i'm going to attempt y=cos(t) and y=sin(t) on interval 0≤x≤pi

OpenStudy (anonymous):

But this question is very much solved- thanks again!

OpenStudy (anonymous):

np, if you get stuck I'll be on for a little bit. BTW a tip for you next question there is no method for finding the integral of trig functions, something you just gotta memorize.

OpenStudy (anonymous):

Sounds good- i'll certainly be reviewing this conversation

OpenStudy (anonymous):

=)